OpenStudy (anonymous):

A savings account compounds interest, at a rate of 17%, once a year. John puts $1,000 in the account as the principal. How can John set up a function to track the amount of money he has?

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\] A = Amount at maturity or at time t. P = Principal Amount = $1,000 r = Annual interest rate in decimal = 0.17 n = compounding period (compounded how many times a year) = 1 t = years invested

OpenStudy (anonymous):

which one is the interest rate?

OpenStudy (ranga):

r is the interest rate = 17% = 0.17 \[\Large A = 1000(1 + \frac{ 0.17 }{ 1 })^{1*t} = 1000(1.17)^t\]